OpenStudy (anonymous):

what is -5(a+4)=15 WILL GIVE MEDAL

OpenStudy (kendricklamar2014):

First step: Simplify both sides to get:\[-5a - 20 = 15\]

OpenStudy (lexi724):

www.cymath.com will save your life lol

OpenStudy (kendricklamar2014):

Now, we add 20 to both sides:\[-5a - 20 + 20 = 15 + 20\] Which equals:\[-5a = 35\]

OpenStudy (kendricklamar2014):

Last step we divide both sides by -5:\[\left(\begin{matrix}-5a \\ -5\end{matrix}\right)\left(\begin{matrix}35 \\ -5\end{matrix}\right)\]

OpenStudy (kendricklamar2014):

Which gives you the answer:\[a = -7\]

OpenStudy (anonymous):

THANKYOU THAT MAKES SENSE!

OpenStudy (kendricklamar2014):

\(\color{blue}{Your ~ Welcome }\)